hi! so I'm almost done with my homework and I'm unsure of how to do this question. the first part is multiple choice and the blank is asking for a solution, thank you!
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Explain how the graph of the function ​f(x)=1/(x-7)^2 can be obtained from the graph of y=1/x^2. Then graph f and give the domain and range.
To obtain the graph of​ f, shift the graph of y=1/x^2 _____ (up/down/to the left/to the right) _____ units
(enter your response here ​unit(s).)

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

the graph of y = 1 / (x-7)^2 is the same graph as the graph of y = 1 / x^2 after it is shifted to the right by 7 units.
here's my graph,....



the graph of y = 1/x^2 is in red.
the graph of y = 1/(x-7)^2 is in blue.

you can see that the blue graph is the same as the red graph, ,only it has been shifted to the right 7 units.

the domain of 1/x^2 is equal to all values of x except at x = 0.
the domain of 1/(x-7)^2 is equal to all values of x except at x = 7
the range of both equations is equal to all values of y greater than 0.

Answer by ikleyn(53937)

.

Hey,

there is free of charge plotting tool in the common access in the Internet

www.desmos.com/calculator


Using it, you can do any plot of any function ON YOUR OWN.

For it, you only need to print the formula of the function in the port window - - - the plotter will do the rest for you momentarily.

So, using this plotter, you can plot the parent function and the "child" function
and see by your own eyes how the child function is obtained from the parent function.